Tim Smunt has been asked to evaluate two machines. After some​ investigation, he determines that they have the costs shown in the following​ table:

                                                                                                  

Machine A

Machine B

Original Cost

$15,000

$24,000

Labor per year

$2,200

$4,400

Maintenance per year

$4,300

$1,200

Salvage value

$1,600

$7,200

He is told to assume​ that:

1. The life of each machine is 3 years

2. The company thinks it knows how to make 14​% on investments no more risky than this one.

3. Labor and maintenance are paid at the end of the year.

The NPV for Machine A = ______________ ​(round your response to the nearest whole number and include a minus sign if​ necessary).
Answer #1

For Machine A -

Cash Flow at Year 0 = -15000

Cash Flow at Year 1 = -2200-4300 = -6500

Cash Flow at Year 2 = -6500

Cash Flow at Year 3 = -6500+1600 = -4900

NPV for each cashflow can be calculated by below mentioned formula -

NPV = Cashflow/(1+i)^n

Where,

i = interest rate

n = number of years

NPV = -15000 - 6500/(1.14) - (6500/(1.14)^2) - (4900/(1.14)^3)

NPV = -15000 - 5701.75 - 5001.54 - 3307.36

NPV = -29010.65

NPV = -29011